Solve by any method. Assume a and b represent nonzero constants.
ax + by = 3
ax - by = 10

Answer by adamchapman(301)

ax+by=3 ..1
ax-by=10 ..2
adding equations (..1) and (..2) together gives:
2ax=13 ..3
ax=13/2=6.5 ..4
substitute equation (..4) into (..1) to give:
6.5+by=3
by=-3.5

I hope this helps.
P.S. I am trying to start my own mathematics homework help website. I would be extremely grateful if you would send me your feedback on my help to adam.chapman@student.manchester.ac.uk.